OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given number as solutions. Type the equation in standard form, ax^2+bx+c=0

OpenStudy (anonymous):

If the given number are say m and n then x-m)(x-n)=0 x^2-(m+n)x+mn=0

Directrix (directrix):

@ Jes ==> What are the given numbers that are solutions?

OpenStudy (anonymous):

If the given solutions are u and v then the equation is : \[x ^{2} - (u+v)x + uv = 0\]